# Memory-augmented Transformers can implement Linear First-Order Optimization Methods

### v1

@@ 3.1 Dynamic Memory for CGD-like Algorithms
where $\gamma_{\ell}$ and $\alpha_{\ell}$ control past update influence and step size
Proof Sketch.  CGD refines search directions using current gradients and previous directions. The Transformer simulates this by using $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ as the current update, analogous to the gradient in CGD, and $\gamma_{\ell}\mathbf{R}_{\ell-1}$ to refine the previous search direction, corresponding to the recursive update of $\mathbf{s}_{n}$ in CGD.
The recursive update for $\mathbf{R}_{\ell}$ thus mimics $\mathbf{s}_{n}$, the search direction in CGD. The update for $\mathbf{Z}_{\ell+1}$ uses $\mathbf{R}_{\ell}$, scaled by $\alpha_{\ell}$, similarly to how CGD iterates are updated using $\mathbf{s}_{n}$. With $\mathbf{A}_{\ell}=\mathbf{I}$, this process matches CGD applied to the loss $R_{\mathbf{w}^{*}}(\mathbf{w})$ (2.11), using both current and previous gradients to refine the search direction. $\square$
3.2 Implementing $k$ Steps of LFOM with Memory Registers
We extend our analysis to show how Transformers can simulate $k$ steps of Linear First-Order Methods (LFOMs). This is achieved by maintaining a memory register at each layer, which stores accumulated updates from previous layers, simulating iterative optimization.
Proposition 2.
A memory-augmented Transformer implements $k$ steps of LFOM by maintaining memory registers across layers, where the update rules are:
@@ 3.2 Implementing $k$ Steps of LFOM with Memory Registers
and where $\Gamma_{j}^{\ell}$ governs the contribution of previous layers, and $\odot$ is the Hadamard product for scaling.
Proof Sketch.  Memformers with this architecture simulate iterative optimization by refreshing the memory register $\mathbf{R}_{\ell}$ at each layer with $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$, capturing the current update. The cumulative update to $\mathbf{Z}_{\ell+1}$ incorporates past layers through a weighted sum of previous memory registers $\mathbf{R}_{j}$, with weights $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$, mimicking LFOM’s cumulative iterative process.
The Hadamard product $\odot$ modulates the influence of $\mathbf{R}_{j}$, analogous to gradient preconditioning. This setup subsumes the case of diagonal preconditioners $\Lambda_{j}^{k}$ acting on gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{j}^{\mathrm{gd}})$, which in the general form looks like:
@@
The matrices $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$ and $\Lambda_{i}^{k}\in\mathbb{R}^{d\times d}$ serve similar roles, but their dimensions differ. We expect this Hadamard product memory architecture to be able to perform richer algorithms than LFOMs, though a formal characterization of its full potential remains to be done.
The full proof follows from the cumulative memory structure and the connection between attention and preconditioned gradients, as discussed in the proof steps of Lemma 2.11. $\square$
Remark. The update in (3.4) can be interpreted as a form of "Gated Memory," analogous to gating mechanisms in LSTMs and GRUs, which also utilize the Hadamard product to regulate information flow through gates. This resemblance suggests that insights from these architectures may help refine memory mechanisms in Transformers, potentially improving their ability to handle long-term dependencies in optimization tasks. However, further investigation is needed to fully understand this connection.
If the $\Lambda_{i}^{k}$ terms in (3.5) are allowed to be general matrices and the product $\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$ is replaced by the Hadamard product $\Lambda_{i}^{k}\odot\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$, we obtain a class of algorithms that sits between first-order and second-order optimization methods. In fact, this is the type of algorithm implemented by the forward pass of the Memformer under the architecture defined in (3.4). A challenging task for future explorations would be to extract a novel high-level optimization algorithm from the trained Transformer/Memformer; doing so, in a rigorous manner remains out of current reach.
3.3 Experimental Results: Memformers implementing LFOMs
@@
In this section, we present our empirical results for Memformers implementing conjugate gradient descent (CGD), general linear first-order methods (LFOMs), and general LFOMs with $\mathrm{GD}^{++}$. The method $\mathrm{GD}^{++}$ is a quasi-Newton method where the inverse Hessian in Newton’s method is approximated by a truncated Neumann series; for more details on $\mathrm{GD}^{++}$, refer to Section A.10 of Von Oswald et al. (2023a).
We consider the in-context loss function (2.11) for linear regression. The input dimension is set to $d=5$, and the number of training observations in the prompt is $n=20$. Both the inputs $\mathbf{x}^{(i)}$ and the target weight vector $\mathbf{w}^{*}$ are sampled from Gaussian distributions: $\mathbf{x}^{(i)}\sim\mathcal{N}(0,\mathbf{\Sigma})$ and $\mathbf{w}^{*}\sim\mathcal{N}(0,\mathbf{\Sigma}^{-1})$, where $\mathbf{\Sigma}=\mathbf{U}^{\top}\mathbf{D}\mathbf{U}$. Here, $\mathbf{U}$ is a uniformly random orthogonal matrix, and $\mathbf{D}$ is a fixed diagonal matrix with entries $\operatorname{diag}(1,1,0.5,0.25,1)$.
We optimize the function $f$ (2.7) for a three-layer linear transformer using the ADAM optimizer. The matrices $\mathbf{A}_{0}$, $\mathbf{A}_{1}$, and $\mathbf{A}_{2}$ (as in (2.8)) are initialized with independent and identically distributed (i.i.d.) Gaussian entries. Each gradient step is computed using a batch of size 1000, and we resample the batch every 100 steps. We clip the gradient of each matrix to have a maximum norm of 0.01. All plots are averaged over five runs, each with a different randomly sampled $\mathbf{U}$ (and thus different $\mathbf{\Sigma}$).
Figure 1 illustrates the implementation of a CGD-like algorithm under the architecture given by (3.2). In Figure 1(a), the line-search parameters $\alpha_{\ell}$ and deflection parameters $\gamma_{\ell}$ for each layer $\ell$ are obtained by training using ADAM. By “CGD-like,” we mean that upon training the memformer using ADAM, the memformer layers learn general parameters $\alpha_{\ell}$ and $\gamma_{\ell}$ which, while they may not match the exact CGD parameters for individual observations, perform well enough on each observation to be comparable to, if not competitive with, CGD. We further explain the important issue of learning general parameters in Section 4.
Figure 1(b) presents the same experiment as Figure 1(a), using the architecture in (3.2), but with the parameters $\mathbf{A}_{\ell}$ for each layer not restricted to scalars. Thus, past gradients are accounted for, similar to CGD, but with preconditioners $\mathbf{A}_{\ell}$. This is therefore not a “CGD-like” algorithm. We aim to demonstrate that once we allow preconditioned gradients, a memformer implements a certain LFOM that distinctly outperforms CGD.
Figure 2 presents the performance of LFOM under the architecture in (3.4), where the matrix parameters $\Gamma_{j}$ for each layer $j$ are obtained by training using ADAM. In our experiments, we consider the special case of $\Gamma_{j}^{\ell}=\Gamma_{j}\ \forall\ell$, which is more natural, if we consider that each layer $j$ of the memformer has an associated $\Gamma_{j}$. Figure 2(a) shows the results on non-isotropic data, and Figure 2(b) shows the results on isotropic data. Note that this algorithm is quite similar in nature to the previous case in Figure 1(b). Here, the $\Gamma_{j}$’s essentially act as preconditioners of the gradients computed in each layer. Consequently, the graphs of Figures 1(b) and 2(a) are nearly identical. In the isotropic data experiment (Figure 2(b)), we observe that the memformer does not perform better than a linear transformer. In quadratics with isotropic data, there is no significant variation in curvature across directions; thus, incorporating past gradients via momentum offers little advantage. Momentum is more beneficial in cases with non-isotropic data.
Figure 3 presents LFOM with $\mathrm{GD}^{++}$ under the architecture in (3.4), where the $\mathbf{B}_{\ell}$ blocks in the $P_{\ell}$ matrices for each layer $\ell$ (2.8) are allowed to be non-zero. Once again, the matrix parameters for each layer $\ell$ are obtained by training using ADAM. In this case, the $\mathbf{B}_{\ell}$ matrices resemble a heavily truncated Neumann series of the inverse $\mathbf{X}\mathbf{X}^{\top}$ (2.12), resulting in a quasi-Newton method. The experiments are conducted on both non-isotropic data (Figure 3(a)) and isotropic data (Figure 3(b)).
(a) Without Preconditioning
(b) With Preconditioning
Figure 1: Comparison of linear transformer and memformer with general CGD-like parameters to actual CGD running separately on each test observation.
(a) Non-Isotropic Data
(b) Isotropic Data
Figure 2: LFOM performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
(a) Non-Isotropic Data
(b) Isotropic Data
Figure 3: LFOM GD++ performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
4 Experiments: Memformer LFOM Performance vs CGD
@@
We emphasize here that the results presented in Section 3.3 compare the performance of Transformers and Memformers (which learn generic parameters) against CGD that runs on fresh observations of batch size $B=1000$, independently resampled from the same distribution. But unlike CGD that computes specific parameters for each observation, the Transformer and Memformer models learn shared parameters $P_{\ell}$, $Q_{\ell}$ (and $\alpha_{\ell}$, $\gamma_{\ell}$, or $\Gamma_{\ell}$) for each layer $\ell$, and these parameters are applied uniformly across all 1000 observations in the batch. In contrast, CGD is executed individually on each of the 1000 observations in the batch, and the average log-loss versus layers is plotted.
@@
The advantage of Memformers in implementing LFOMs (with matrices $\Gamma_{\ell}$ restricted to scalar multiples of the identity) becomes even more pronounced when tested on training data with small batch sizes, such as $B=1$ and $B=10$. In these scenarios, the Memformers learn parameters that significantly outperform CGD running in parallel on each of the observations in those small batches. Figure 4 demonstrates this comparison.
We further provide an experimental comparison of Memformer implementing LFOM vs. Nesterov Accelerated Gradient Method and Momentum GD in the Appendix.
(a) Batch Size $B=1$
(b) Batch Size $B=10$
@@
Figure 4: Memformer with scalar preconditioners vs. CGD performance on small batch training data ($B=1$ and $B=10$). The memformer demonstrates superior performance on the training data.
5 Experiments: Impact of Using Multi-Headed Attention
@@
Our experiments demonstrate that increasing the number of attention heads leads to better test loss performance. Transformers with multiple attention heads improve test loss by learning diverse preconditioning matrices, allowing them to better adapt to varying data covariance structures. In our architecture, as described in (3.1), the attention values from each head are summed before being assigned to the memory register $\mathbf{R}_{\ell}$ at each layer $\ell$.
Each head captures different aspects of the data, estimating gradients from multiple perspectives. This results in faster convergence, as the heads function like an ensemble, combining different gradient estimates to improve optimization. Additionally, combining gradient estimates from multiple heads acts as implicit regularization by smoothing the gradient updates, stabilizing the learning process, and preventing overfitting. This leads to better generalization on test data.
Figure 5 compares the performance of models with 1-head and 5-head attention. The results illustrate the benefits of using multiple attention heads on convergence speed and test loss performance.
(a) 1-Head Attention Performance
(b) 5-Head Attention Performance
@@
Figure 5: Comparison of memformer (with scalar preconditioners) performance using 1-head and 5-head attention, relative to CGD.
6 Discussion and Future Work
This work demonstrates the capability of memory-augmented Transformers (Memformers) to implement a broad range of first-order optimization methods, opening several research directions. We briefly comment on some of these aspects below.
  1. (i)
    Architectural Flexibility: Small modifications, such as (gated) memory registers, significantly enhance Transformers’ ability to learn and implement diverse optimization algorithms. Future research could explore further architectural innovations to unlock even greater capabilities.
  2. (ii)
    General Function Classes: While our approach successfully makes Transformers implement LFOMs on quadratic functions, future work should extend this to more general objective functions. Doing so may require novel training strategies, and possibly architectural adjustments to handle non-quadratic functions. The role of nonlinear attention and the MLP component of Transformers may also prove to be useful here.
  3. (iii)
    Efficiency vs. Generalization: Attention-based methods require more computation than directly implementing conjugate gradient descent or momentum GD. However, Transformers excel in learning general parameters, enabling LFOMs to generalize across new data without needing per-instance optimization. Exploring practical use of such “learned optimizers” to either warmstart a solver, or to potentially even bypass it, is a tantalizing research topic.
  4. (iv)
    Theoretical Foundations and Convergence Analysis: Strengthening the theoretical basis of Transformers’ optimization capabilities, including convergence analysis and their alignment with classical optimization theory, is another important direction for future research.
  5. (v)
    Meta-learning and Transfer Learning: The ability of Transformers to learn and generalize optimization algorithms offers exciting potential for meta-learning and transfer learning, providing new opportunities in areas where traditional optimization methods fall short.
6.1 Limitations
We briefly remark on some limitations of our current framework. For instance, while LFOMs are quite versatile, our experiments (Figure 1) indicate they do not radically outperform preconditioned GD on general quadratic problems as in (2.11), where the preconditioner matrix $\Gamma_{\ell}$ (and likewise, $\mathbf{A}_{\ell}$) for the current layer $\ell$ is the main contributor to loss performance at each update step $\ell$ (3.1). On the other hand, this behavior is likely due to the task being quadratic, and a future study that tackles more general ICL formulations will likely shed light here.
Transformers can implement second-order methods like Newton’s method (Fu et al., 2023; Giannou et al., 2024), which typically outperform LFOMs in convergence speed and accuracy. However, we reiterate that the main focus of our paper is to explore the space of first-order optimization algorithms that augmented Transformers can learn, as opposed to looking for “the best” algorithm.
7 Reproducibility Statement

### v3

@@ 3.1 Single Dynamic Memory for CGD-like Algorithms
and $\gamma_{\ell},\alpha_{\ell}$ control the influence of past updates and step size, respectively.
Proof Sketch.  Here, $\mathbf{R}_{\ell}$ tracks a single memory register across layers. CGD updates its search direction by combining the current gradient with the previous direction (2.13). The Transformer mimics this by combining $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ (analogous to the gradient) and $\gamma_{\ell}\mathbf{R}_{\ell-1}$, and updating the next state $\mathbf{Z}_{\ell+1}$ via (3.2). With $\mathbf{A}_{\ell}=\mathbf{I}$, this matches CGD applied to (2.11); see Appendix A for details. $\square$
3.2 Implementing $k$ Steps of LFOM with Separate Memory Registers at Each Layer $\ell$
Next, we show how Transformers can simulate $k$ steps of general LFOM updates (1.1) by maintaining a separate memory register at each layer $\ell$.
Theorem 3.2.
A memory-augmented Transformer can implement $k$ steps of LFOM in its forward pass by maintaining separate memory registers across layers $\ell$:
@@ 3.2 Implementing $k$ Steps of LFOM with Separate Memory Registers at Each Layer $\ell$
where $\Gamma_{j}^{\ell}$ weights previous layer updates and $\odot$ is a Hadamard product for scaling.  [^1] The update (3.4) can be interpreted as a form of gated memory. Analogies with LSTMs or GRUs (which also use Hadamard gating) suggest potential ways to refine memory usage in Transformers.
Proof Sketch.  Here, $\mathbf{R}_{\ell}$ is a separate register at each layer. At layer $\ell$, $\mathbf{R}_{\ell}$ captures the current update $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$. The final output $\mathbf{Z}_{\ell+1}$ is updated by summing over registers $\mathbf{R}_{j}$ scaled through $\Gamma_{j}^{\ell}\in\mathbb{R}^{(n+1)\times(d+1)}$. This resembles cumulative gradient steps of general LFOMs with diagonal preconditioners (1.1). A full proof is in Appendix A. $\square$
The Hadamard product $\odot$ modulates the influence of $\mathbf{R}_{j}$, analogous to gradient preconditioning. This setup subsumes the case of diagonal preconditioners $\Lambda_{i}^{k}$ acting on gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$, which in the general form looks like:
@@
The matrices $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$ and $\Lambda_{i}^{k}\in\mathbb{R}^{d\times d}$ serve similar roles, but their dimensions differ. We expect this memory architecture to be able to perform richer algorithms than LFOMs, though a formal characterization of its full potential remains to be done.
Remark 3.3.
Crucially, unlike classical methods such as CGD, which “adapt” by computing step sizes ($\alpha$) and conjugacy coefficients ($\gamma$) separately for each sample—even within the same batch—our Memformer uses a single set of learned parameters shared across all samples. For a test batch of size $B$, CGD thus effectively uses $2B$ parameters per iteration, whereas the Memformer has just the trained internal parameters $\{P_{\ell},Q_{\ell},\Gamma_{\ell}\}$ (3.4). This amounts to a meta-optimizer that can exploit “average-case” behavior over the data distribution, rather than re-optimizing each instance individually.
3.3 Experimental Results
@@
We evaluate Memformers for learning CGD, general LFOMs, and LFOMs with $\mathrm{GD}^{++}$ (a quasi-Newton method approximating the Hessian inverse via a truncated Neumann series (Von Oswald et al., 2023a).
Setup. We use the in-context loss (2.11) for linear regression with $d=5$ and $n=20$. Each run samples a random orthogonal matrix $\mathbf{U}\in\mathbb{R}^{d\times d}$ and uses $\mathbf{D}=\mathrm{diag}(1,1,\tfrac{1}{2},\tfrac{1}{4},1)$. With a variance scalar $\sigma^{2}$ (default $1$), we set $\mathbf{\Sigma}=\sigma^{2}\mathbf{U}^{\top}\mathbf{D}\mathbf{U}$. We draw inputs $\mathbf{x}^{(i)}$ i.i.d. from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$, and targets $\mathbf{w}^{*}$ from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}^{-1})$. We optimize the in-context prediction function $f$ (2.7) using a 4-layer linear Transformer trained via ADAM, initializing parameters $\mathbf{A}_{0},\mathbf{A}_{1},\mathbf{A}_{2}$, $\mathbf{A}_{3}$ (2.8) i.i.d. Gaussian, with batch size $1000$ (resampled every 100 steps) and gradient clipping of $0.01$. Plots average over five independent runs with different $\mathbf{U}$.
CGD-like Architecture. Figure 1 shows a Memformer parameterized as in (3.2), where the globally learned $\alpha_{\ell}$ and $\gamma_{\ell}$ emulate CGD’s line search and deflection. Unlike standard CGD, which refines these parameters individually for each sample, our Memformer uses a single set of learned scalars shared across all data. Despite this approximation, performance remains competitive. Figure 2 extends the approach by allowing non-scalar preconditioners $\mathbf{A}_{\ell}$ (2.8), thus making the method a more general “LFOM-like” strategy that can even outperform standard CGD.
LFOM Memformer. Figure 3 shows a Memformer with the LFOM-like architecture (3.4), where each layer’s $\Gamma_{\ell}$ acts as a preconditioner. In our experiments, we consider the special case of $\Gamma_{j}^{\ell}=\Gamma_{j}\ \forall\ell$, which is more natural, if we consider that each layer $j$ of the Memformer has an associated $\Gamma_{j}$. In practice, these parameters further improve performance. Figure 4 extends LFOM with $\mathrm{GD}^{++}$ by allowing the $\mathbf{B}_{\ell}$ blocks (2.8) to be non-zero. In this case, the $\mathbf{B}_{\ell}$ matrices resemble a heavily truncated Neumann series of the inverse $\mathbf{X}\mathbf{X}^{\top}$ (Hessian of (2.11)). Our experiments with more than 4 layers are detailed in Appendix C.
Figure 4: LFOM Memformer with GD++ (3.4) vs. CGD, where the $\mathbf{B}_{\ell}$ blocks (2.8) approximate the Hessian inverse (quasi-Newton).
3.4 Influence of Batch Size on Performance
@@
The results in Section 3.3 compares the performance of Transformers and Memformers (which learn shared generic parameters during training) against CGD, which computes specific parameters for each observation in a batch of size $B=1000$, independently resampled from the same distribution. While CGD optimizes each observation individually, Transformers and Memformers use shared parameters $P_{\ell},Q_{\ell}$ (and $\alpha_{\ell},\gamma_{\ell}$, or $\Gamma_{\ell}$) for each layer $\ell$, uniformly used for all 1000 observations in the batch. The average log-loss versus layers are plotted for comparison in Figures 2 and 3.
@@
The strength of LFOM Memformers (3.4) (with matrices $\Gamma_{\ell}$ restricted to scalar multiples of the identity) becomes even more pronounced when trained on smaller batch sizes, such as $B=1$ and $B=10$. In these cases, Memformers learn parameters that significantly outperform CGD, even when CGD runs independently on each observation in the batch. Figures 5 and 6 illustrate this comparison.
@@
Figure 5: LFOM Memformer (3.4) with scalar preconditioners $\Gamma_{\ell}$ vs. CGD performance on small batch size ($B=1$). The Memformer demonstrates superior performance.
Figure 6: LFOM Memformer (3.4) with scalar preconditioners $\Gamma_{\ell}$ vs. CGD performance on small batch size ($B=10$). The Memformer outperforms CGD on the training data.
4 Multi-Headed Attention and Out-of-Distribution (OOD) Performance
4.1 OOD Adaptation using Multi-Headed Attention
@@
Next, we explore how multi-head attention further improves both in-distribution and OOD performance. In our experiments, increasing the number of heads yields notable gains in test-loss metrics. By learning diverse preconditioning matrices, multi-head architectures help Memformers adapt to data with varying covariance structures. Concretely, in the update (3.1), we sum the attention outputs from different heads into a single memory register $\mathbf{R}_{\ell}$ at each layer, thus aggregating multiple “views” of the underlying data distribution.
@@
Figure 7: Performance of an LFOM Memformer (3.4) (with scalar preconditioners $\Gamma_{\ell}$) comparing one-head and five-head attention, relative to CGD. Multi-head attention accelerates convergence and improves test performance.
These findings align with prior studies underscoring the critical role of multi-head attention in in-context learning. For example, Chen et al. (2024) highlight its importance in sparse linear regression, and Cui et al. (2024) demonstrate its theoretical and empirical benefits. As illustrated in Figure 7, the number of heads strongly influences both convergence speed and final test loss.
4.2 Test-Time Adaptation and Gated Memory
Building on the above, we next consider test-time adaptation and show how gating mechanisms allow robust performance under distribution shift.
Mixture of Experts (MoE) for Variance Adaptation. To extend multi-head architectures to OOD scenarios, we employ a mixture of experts (MoE) approach Jacobs et al. (1991) in a 3-head, 4-layer Memformer. Each head specializes in Gaussian inputs with different variances $\sigma^{2}\in\{1.0,2.0,3.0\}$. At inference, a learnable gating mechanism scales the layerwise output $\sum_{j=0}^{\ell}\Gamma_{j}\odot\mathbf{R}_{j}$ (3.4) per head coordinate-wise. Concretely, each head $H_{i}$ has a scalar gating coefficient $\alpha_{i}\in\mathbb{R}$, enabling the model to adapt effectively to unseen variances (e.g., $\sigma^{2}=\{0.5,1.5,3.5\}$). As shown in Figure 8, this design ensures robust generalization even when the variance differs from the training set.
Figure 8: MoE-based 3-head, 4-layer Memformer (3.4) evaluated on Gaussian inputs of varying unseen variances. A trainable gating mechanism allows effective OOD adaptation.
Gaussian Mixture Models (GMMs). We further examine OOD adaptation in a 3-head, 4-layer Memformer tested on a Gaussian Mixture Model (GMM) with different means ($\{2.5,0.0,-1.0\}$), variances ($\{0.5,1.0,2.0\}$), and mixture weights ($\{0.2,0.3,0.5\}$). Each head specializes in one mixture component. As before, a gating mechanism similarly scales the layerwise output $\sum_{j=0}^{\ell}\Gamma_{j}\odot\mathbf{R}_{j}$ across heads.
For this GMM setup, we observe that using a 3-head, 4-layer Linear Transformer (cf. Figure 9) gives better performance than a 3-head, 4-layer Memformer (cf. Figure 10), as the $\Gamma_{j}$ matrices of each head in the Memformer overfit to the individual GMM components. However, this issue can be overcome, with a higher test-time budget, if the $\Gamma_{j}$ matrices are also allowed to be trainable at test time (see Figure 10).
Figure 9: A 3-head, 4-layer Linear Transformer tested on a GMM. Each head specializes in one mixture component, and a gating module adjusts the model for OOD data.
Figure 10: A 3-head, 4-layer Memformer (3.4) tested on a GMM. The gray line shows results with a learned gating scheme, while the black line also includes test-time updates to the $\Gamma_{j}$ matrices.
Theorem 4.1 (Multi-Head Memformer with Soft Gating).
Consider a multi-head Memformer with $H$ heads, each parameterized by $\{P_{\ell}^{h},Q_{\ell}^{h},\Gamma_{\ell}^{h}\}$. Suppose $\mathbf{Z}_{0}$ is drawn from a mixture of $M$ Gaussian components $\{\mathcal{N}(\mu_{m},\Sigma_{m})\}_{m=1}^{M}$, where each $\mathbf{x}^{(i)}\sim\mathcal{N}(\mu_{m},\Sigma_{m})$ and $\mathbf{w}^{*}\sim\mathcal{N}(\mu_{m},\Sigma_{m}^{-1})$. If each component $m$ has at least one head $h_{m}$ achieving near-optimal first-order performance, then there exist scalar gating coefficients $\alpha_{m,h}$ such that for any new prompt from component $m$, the Memformer achieves near-optimal performance via soft-gating:
$\alpha_{m,1}\,\mathrm{Head}_{1}\;+\;\cdots\;+\;\alpha_{m,H}\,\mathrm{Head}_{H}.$
[…]
A formal proof of Theorem 4.1 appears in Appendix A. As indicated above, a Transformer can handle OOD data by fine-tuning specific test-time parameters (e.g., the gating coefficients or the diagonal of $\Gamma_{i}$, as in Figure 10). This design in effect operates as a meta-version of in-context learning: the main architecture parameters remain fixed, while gating and memory preconditioners can adjust to accommodate distribution shifts, depending on the test-time budget.
5 LFOMs as Learnable Algorithms
Finally, LFOM updates can themselves be learned as a global meta-optimizer, bridging the gap between classical algorithms and in-context algorithmic learning. One can directly learn the parameters of a linear first-order method (LFOM) from data, treating it as a global algorithm that generalizes across random in-context samples. The following theorem provides a standard finite-sample guarantee for such an approach under mild regularity conditions (boundedness, sub-Gaussianity, continuity, and compactness). A complete version and proof appear in Appendix A.
Theorem 5.1 (Statistical Learnability of LFOMs in the In-Context Setting).
Fix a class of LFOMs parameterized by $\theta\in\Theta\subset\mathbb{R}^{p}$, e.g., via diagonal preconditioners. Each LFOM $\theta$ maps an in-context prompt $\mathbf{Z}_{0}$ to a final prediction $\hat{y}_{\theta}(\mathbf{Z}_{0})$ for $\mathbf{x}(n+1)$, incurring the loss $\ell(\theta;\mathbf{Z}_{0})=\bigl{(}\hat{y}_{\theta}(\mathbf{Z}_{0})-\langle\mathbf{x}(n+1),\mathbf{w}^{*}\rangle\bigr{)}^{2}$.
Drawing $N$ i.i.d. prompts $\{\mathbf{Z}_{0}^{(i)}\}_{i=1}^{N}$ from $\mathcal{D}$, define $\widehat{\theta}=\arg\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\ell\bigl{(}\theta;\mathbf{Z}_{0}^{(i)}\bigr{)}$. Then, with high probability over the sample draw,
$\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\widehat{\theta};\,\mathbf{Z}_{0})\bigr{]}\;\leq\;\min_{\theta\in\Theta}\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\theta;\,\mathbf{Z}_{0})\bigr{]}\;+\;\epsilon(N,\delta),$
where $\epsilon(N,\delta)\to 0$ as $N\to\infty$. In other words, the learned LFOM parameters $\widehat{\theta}$ generalize well on new in-context samples from $\mathcal{D}$.
Note that our Memformer update rule (3.4) encompasses an even broader family of optimization strategies than classical LFOMs. Figure 11 illustrates how learning LFOM parameters (e.g. diagonal preconditioners) directly from data (using ADAM or similar) can match or exceed the performance of Conjugate Gradient Descent (CGD), while using one shared set of parameters across multiple samples. Such a strategy can overfit if batches are extremely small, but tends to generalize effectively for moderate to large batch sizes.
Figure 11: Comparison of a learned LFOM approach vs. standard Conjugate Gradient Descent (CGD) on a synthetic quadratic problem (2.11). LFOM parameters (3.5) are trained directly from data, while CGD is run separately on each instance. ($B$ = Batch Size)
6 Conclusion and Future Work
We find that memory-augmented Transformers (Memformers) can learn and implement a wide range of first-order optimization algorithms, including gradient descent, conjugate gradient, and momentum methods. This versatility highlights their potential as general-purpose meta-optimizers. Below are several avenues for further exploration:
Architectural Extensions, MoE, and TTA. Additional gating strategies, mixture-of-experts (MoE), and refined memory modules may boost performance on out-of-distribution (OOD) data. Even a small number of experts can deliver robust gains, though identifying the “right” number of experts remains open. More advanced test-time adaptation (TTA), such as fine-tuning the gating or expert parameters on new data, could further enhance generalization to shifting distributions.
Beyond Quadratic Objectives. Most theoretical work on algorithm learning in Transformers has focused on linear regression, leaving open the understanding of more complex, potentially nonconvex objectives. By integrating nonlinear attention layers or MLP components, Memformers could tackle tasks like logistic regression or PDE-based inverse problems, where simple quadratic assumptions are insufficient. Analyzing Memformers’ in-context learning on such domains could unify meta-learning research with more realistic, large-scale optimization problems, while clarifying theoretical guarantees beyond the linear-quadratic case.
Efficiency vs. Generalization. While attention-based models can be more computationally demanding than methods like momentum or CGD, they exhibit strong generalization across diverse tasks. Systematic studies of runtime and accuracy trade-offs could clarify the conditions under which learned optimizers outperform classical approaches.
Meta-Learning and Transfer. Memformers naturally connect to meta-learning by reusing learned optimization strategies. Future investigations might examine more sophisticated TTA, gating, or lightweight fine-tuning approaches to handle distribution shifts. Understanding how best to structure MoE—such as how many experts to deploy and how to select them on the fly—promises both practical efficiency and robust transfer in a variety of domains.
Limitations of our framework.
While Memformers are versatile, our experiments (Figures 2, 3) show they do not drastically outperform advanced methods on quadratic tasks (2.11), likely due to the task structure. Future work on broader ICL formulations may shed light on this matter. Notably, Transformers can implement second-order methods like Newton’s (Fu et al., 2023; Giannou et al., 2024), which typically converge faster and more accurately than LFOMs. However, our focus in this paper is on exploring first-order optimization algorithms that augmented Transformers can learn, rather than proposing a one-size-fits-all optimizer.
Reproducibility Statement